OpenStudy (anonymous):

if Bc=15 and DC=28 what is the diameter of the circle O to the nearest tenth A.13.1 B.37.3 C.52.3 D.67.3

OpenStudy (anonymous):



OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

have a look at this page http://www.regentsprep.org/Regents/math/geometry/GP14/CircleSegments.htm and use theorem 3

OpenStudy (anonymous):

i still dont understand what to do though ?

jimthompson5910 (jim_thompson5910):

let AB = x

jimthompson5910 (jim_thompson5910):

we want to find the length of AB, ie solve for x

jimthompson5910 (jim_thompson5910):

using theorem 3, we would get BC*AC = (DC)^2 15*(15+x) = 28^2 x = ???

OpenStudy (anonymous):

TY

jimthompson5910 (jim_thompson5910):

what did you get?

OpenStudy (anonymous):

52.3

jimthompson5910 (jim_thompson5910):

Try it again. That is incorrect.

OpenStudy (anonymous):

how is that incorrect , i work the problem out and got 52.2 rounded to the nearest tenth it would be 52.3

jimthompson5910 (jim_thompson5910):

I'll show you how to solve for x 15*(15+x) = 28^2 15*(15+x) = 784 15*15+15*x = 784 225 + 15x = 784 15x = 784 - 225 15x = 559 x = 559/15 x = 37.2666666666667 which rounds to 37.3

OpenStudy (anonymous):

question why would you put 15 twice if BC=15 then formula should be 15(x)=28^2 which would give you 15x=784. Not really understanding how you got another 15.

jimthompson5910 (jim_thompson5910):

if we use theorem 3, we will get this BC*AC = (DC)^2

jimthompson5910 (jim_thompson5910):

AC is composed of the segments AB and BC so AC = AB + BC AC = x + 15 AC = 15 + x

jimthompson5910 (jim_thompson5910):

that's how I'm getting 15 + x instead of just "x"

jimthompson5910 (jim_thompson5910):

notice how they state "(whole secant)×(external part) = (tangent)^2"

jimthompson5910 (jim_thompson5910):

The AC is the whole secant